### v13

@@ 3 The inter-period transfer mechanism for each individual
$\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
for given realised $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
[…]
and
[…]
From the last equality we have for $n\in\mathbb{N}$ $\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e. $[P]$, and $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale as claimed.
Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathrm{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere. ∎
As the proof of Proposition 2 shows, mechanism $\mathcal{M}$ functions like a fair-value gamble on lifetime utilities across periods. The lifetime utility in each period is a random variable whose sequence forms a martingale. Consequently, the mechanism is fair in the sense that each individual’s expected lifetime utility remains unchanged from the time of the bet.
4 The inter-period transfer mechanism for the government
However, the convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism. This is because the government predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma. Let $\overline{\tau}_{n}$ be the average transfer to those who were born $n$ periods ago. Then $\overline{\tau}_{n}$ is given by as follows.
$\overline{\tau}_{n}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
Lemma 2.
The average transfer for each age of individuals $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
Proof.
We first consider the case $n=1$. For $n=1$, since $\mathrm{E}[\overline{v}_{1}(\lambda_{1}(e_{1}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathrm{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathrm{E}[e_{0}]=\overline{\tau}_{0}.$
For $n>1$ since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
$\overline{\tau}_{n-1}$ $=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathrm{E}[e_{n-1}]$
$=\mathrm{E}[u^{-1}(\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathrm{E}[e_{n-1}]$
$\leq\mathrm{E}[\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathrm{E}[e_{n-1}]$
$=\mathrm{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathrm{E}[e_{n}]$
$=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
Let $g(t,\lambda_{0})$ be the expected balance of payments in period $t$ for the government for given initial utility control variable $\lambda_{0}$. Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of contimuum of individuals on the $(1-\alpha)$ interval is born in each period, $g(t,\lambda_{0})$ is given by
$g(t,\lambda_{0})$ $=\alpha^{t}(-\overline{\tau}_{t})+\alpha^{t-1}(1-\alpha)(-\overline{\tau}_{t-1})+\cdots+\alpha(1-\alpha)(-\overline{\tau}_{1})+(1-\alpha)(-\overline{\tau}_{0})$
$=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
For the $\mathcal{M}$ mechanism to be sustainable, the balance of payments in each period must be greater than or equal to be zero. As we see in the following proposition, with the approximation of the second order Taylor polynomial, if incomes have a symmetric distribution, a sufficient condition for the balance of payments in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
Proposition 3.
Assume that (i) incomes have a symmetric distribution, (ii) $-u^{\prime\prime}/u^{\prime}$ is a decreasing function, (iii) $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function or equivalently,
$\frac{\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{{}^{\prime\prime}}({\scriptstyle u^{-1}(y)})}{-\left(-\frac{u^{\prime\prime}}{u^{\prime}}\right)^{\prime}({\scriptstyle u^{-1}(y)})}\leq\frac{(u^{-1})^{\prime\prime}(y)}{(u^{-1})^{\prime}(y)},$
and (iv) $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. Then $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ with the approximation of the second order Taylor polynomial.
The proof of Proposition 3 uses the following lemma.
Lemma 3.
Under the assumptions (i), (ii), (iii) in Proposition 3, the second order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
Proof.
@@ 4 The inter-period transfer mechanism for the government
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (6). Using the third order Taylor polynomial, for a given utility control variable $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=$ $u^{-1}(\overline{v}_{1}(\lambda_{n-1}))+\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
@@
$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (10)
where $\Delta_{n-1}=(1-\beta)(e-\mathrm{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$. Taking the expectation of (10), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (10) will cancel out when the expectation is taken.
@@
$\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathrm{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
$\geq\mathrm{E}\left[\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
$=\mathrm{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
@@
We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (7). We change $\Delta_{n-1}$ in (10) to
$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathrm{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (10) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathrm{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
By a similar argument, the second order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
We prove Proposition 3 below.
Proof.
From Lemma 3, with the second order polynomial approximation we have
@@
Hence the following inequality holds.
$g(t,\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
@@
$-\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
and we see a sufficient condition for $g(t,\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Proposition 3 is given by the third order polynomial approximation as follows.
Corollary 1.
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Proposition 3 holds with the second order polynomial approximation if the initial utility control variable $\lambda_{0}$ satisfies
where $\lambda_{E}=1/(u^{\prime}(u(\mathrm{E}[e_{t}])))$.
Proof.
Let us first examine the case where $\lambda_{1}$ can be defined in (6). Applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (12)
On the other hand, applying Tayler theorem to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
@@
$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (13)
where $\Delta_{E}=\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0})$. Since the first term on the right-hand side of (13) is positive, it is sufficient to show the second term on the right-hand side of (13) is equal to or greater than the first term on the righ-hand side of (12). However, from the corollary assumption (11), we have
@@
which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second order polynomial approximation. For the case where $\lambda_{1}$ must be defined in (7), we have
@@
Figure 1: Slope of Utility Function
{singlespace*}
Notes: This figure shows a graph of the slope of a utility function. The horizontal axis is consumption, and the vertical axis is utility. The curve is the utility function. The length $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ is a linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
The left-hand side of the inequality (11) is a linear approximation of the difference between the consumption level that ensures the expected income and the consumption level in the mechanism $\mathcal{M}$ with the utility control variable $\lambda_{0}$. The inequality (11) means that this length is greater than the square root of the weighted income variance. Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$. The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ could be set higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (11).
A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$. If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (11) may be unnecessarily strict.
5 Numerical examples
According to Corollary 1, in the following case, the mechanism $\mathcal{M}$ is sustainable and the lifetime utilities of all generations are higher than that in the state of autarky. Suppose that the utility function is given by
$u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma},$
where $\gamma=-3$. The income set $E$ is $E=\{0,1,2,3,4,5,6,7,8,9\}$. We will refer to this income set as $E_{s}$. The income shocks follow the distribution $P(e_{t}=0)=Pr(e_{t}=9)=0.025$, $P(e_{t}=1)=P(e_{t}=8)=0.045781$, $P(e_{t}=2)=P(e_{t}=7)=0.092762$, $P(e_{t}=3)=P(e_{t}=6)=0.148524$, $P(e_{t}=4)=Pr(e_{t}=5)=0.187933$. We will refer to this distribution as $P_{s}$ Suppose that the discount factor and the probability of being alive are $r=\alpha=0.93$. Then Proposition 3 assumptions are satisfied for $\lambda_{0}\in[\lambda_{CE},133]$, where $\lambda_{CE}=49.608=1/u^{\prime}(E[u(e_{t})])$. Note that $\lambda_{E}=166.375$, the risk premium is $1.826$, and $\sqrt{(1-\alpha r)^{2}/(\alpha r^{2}(1-\alpha))\mathrm{Var}[e_{t}]}$ is $1.147$.
Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals and corresponding transfer balances for $100$ periods for the above case. The initial utility control variable is set $\lambda_{0}=50$.
As Proposition 2 implies, the series of sample means of lifetime utilities does not approach negative infinity, which has been called immiserization. However, as Lemma 2 shows, the corresponding series of sample means of transfer balances $\{-\tau_{n}\}_{n=0}^{100}$ is decreasing.
Figure 2: A series of sample means of lifetime utility and corresponding transfer balance
{singlespace*}
Notes: This figure shows a series of sample means of lifetime utilities and corresponding transfer balances for $100$ periods, where $1000$ individuals are assumed to have the utility $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$. The parameters are set so that the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated using the discrete probability distribution $P_{s}$.
Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. In this society, each individual will be alive in the next period with probalibility $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$. As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$ and the discount factor $r=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated with the same symmetric distribution $P_{s}$.
The series of sample means of transfer balances $\{g(n,\lambda_{0})\}_{n=0}^{1000}$ remains non-negative, and supports Proposition 3.
Figure 3: A series of sample means of balance of payments $\{g(t,\lambda_{0})\}$
{singlespace*}
Notes: This figure shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. As in Figure 2, each individual’s utility is $(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial utility control variable $\lambda_{0}=50$, the discount factor $r=0.93$, the probability of being alive $\alpha=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated using the discrete probability distribution $P_{s}$.
Immiserization observed in previous studies [e.g. Green, 1987, Thomas and Worrall, 1990] is due to the approach to allocate risks within each period across the polulation of individuals. To ensure truthful reporting, only individuals with the highest incomes keep their lifetime utility level in the next period, and lifetime utilities of other individulas in the next period are discounted according to reported incomes. As a result, the average contract value of individuals becomes a decreasing series.
However, the mechanism $\mathcal{M}$ simply shifts risks to future periods. A similar limitation is that, the proposed mechanism leads to a dicreasing series of the average transfer balances for each generation, which could make the government’s balance of payments negative infinity. However, this could be solved by setting the initial utility control variable appropriately, so that average transfers are negative for the first few periods from birth. If this sense of intergenerational cooperation is properly organised, efficient income allocations could be sustainable even under asymmetric information about income shocks.
The property of the mechanism $\mathcal{M}$ contrasts with the numerical results of Marcet and Marimon [1992], where the lifetime utility of agent 1 (manager) is higher in the state of autarky than in the state of the incentive contract with agent 2 (investor). This is probably because, first, Marcet and Marimon [1992] is a growth model with the assumption that capital and the manager’s consumption, increases when the manager chooses the optimal level of investment. Second, since the transfers in the numerical example for incentive contracts in Marcet and Marimon [1992] are not competitive, the external financing contracts end up reducing agent 1’s allocation.
6 Checking the mecanism optimality for each generation
Lastly, we express the mechanism $\mathcal{M}$ using a Bellman equation. We focus on the case that (9) in Lemma 1. First we guess the value function $V:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\to\mathbb{R}\cup\{-\infty\}$ is given by for $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$
$V(\lambda_{t},t)=\mathrm{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}.$ (14)
Second, we verify whether this function satisfies the Bellman equation:
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t},t)}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta V(\lambda_{t+1},t+1),$ (15)
where $\Lambda:\mathbb{R}_{++}\times\mathbb{N}\cup\{0\}\multimap\mathbb{U}$ is an available control correspondence defined such that
$\Lambda(\lambda_{t},t)=\arg\max_{\tau}\left[\lambda_{t}u(e_{t}+\tau)-\tau\ \text{ subject to }(1-\beta)\mathrm{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k}\tau(e_{k})\right]\leq\text{Const.}\right]$. ]
$\mathbb{U}$ is the set of control parameters, and $\lambda_{t+1}$ in (15) is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{U}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (7). We observe the following proposition.
Proposition 4.
If the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ satisfies (9), then $V$ fullfills the Bellman equation and it is the value function of the dynamic optmisation problem, where the control function $\tau^{*}$ maximises $J$ in (2).
A sufficient condition for $V$ to be the value function of the Bellman equation is given by Wiszniewska-Matyszkiel [2011] which is to satisfy the Bellman equation and the following terminal condition.
Terminal condition [Wiszniewska-Matyszkiel, 2011]
  • (i)
    For every $\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$
    $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}\leq 0$
    and
  • (ii)
    for every $\lambda_{t}\in\{\lambda\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$, if $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}<0$, then
    $J(\lambda_{t},t,\hat{\tau})=-\infty$
    for every $\hat{\tau}:\mathbb{R}_{++}\to\mathbb{R}_{+}$ such that $\lambda_{t}$ in $\mathcal{M}$ corresponding to $\hat{\tau}$ and $t$.
Proof.

### v15

@@ Appendix C Proof of Theorem 3
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5). Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))-u^{-1}(\overline{v}_{1}(\lambda_{n-1}))=$ $\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
@@
$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (C.1)
where $\Delta_{n-1}=(1-\beta)(e-\mathbb{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$. Taking the expectation of (C.1), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (C.1) will cancel out when the expectation is taken.
@@
$\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
$\geq\mathbb{E}\left[\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
$=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
@@
We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (6). We change $\Delta_{n-1}$ in (C.1) to
$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathbb{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (C.1) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
By a similar argument, the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
We prove Theorem 3 below.
Proof of Theorem 3.
From Lemma C.1, with the second-order polynomial approximation, we have
@@
Hence, the following inequality holds.
$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
@@
$\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
and we see a sufficient condition for $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
Appendix D Proof of Corollary 1
Proof of Corollary 1.
Let us first examine the case where $\lambda_{1}$ can be defined in (5). Applying Taylor theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (D.1)
On the other hand, applying Taylor to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
@@ Appendix D Proof of Corollary 1
$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (D.2)
where $\Delta_{E}=\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0})$. Since the first term on the right-hand side of (D.2) is positive, it is sufficient to show the second term on the right-hand side of (D.2) is equal to or greater than the first term on the right-hand side of (D.1). However, from the corollary assumption (9), we have
@@
which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second-order polynomial approximation. For the case where $\lambda_{1}$ must be defined in (6), we have
@@
Appendix E Proof of Proposition 1
Proof of Proposition 1.